Ask
your own question, for FREE!
Probability
14 Online
A coin is weighted so that the probability of getting heads is 0.70. If the coin is flipped twice, what is the probability of getting a tail followed by a head?
Still Need Help?
Join the QuestionCove community and study together with friends!
P(heads) = 0.70 P(tails) = 0.30 P(tails followed by heads) = 0.30 * 0.70
where does .30 come from
The total probability is 1. If the probability of heads is 0.70, and the onty other possible outcome is tails, that means the probability of tails is 1 - 0.7 = 0.3
thank you
makes sense now
Still Need Help?
Join the QuestionCove community and study together with friends!
You're welcome.
wait so the answer is 0.21???
@mathstudent55 @cabette ^^
Yes.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
3 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal